Name X-Value Y-Value 4 (a) Plot the five points and the centroid. Be sure that Y is the vertical axis and X is the horizontal axis. (b) Draw a line that appears to fit the five points. (c) Find the bo and b, for the line you drew in part (b). (d) Find the residuals. Specifically, calculate e,Y (bo bX) for i H J, K, L, and M (e) Calculate the sum of the squared residuals (2 e) for the line drawn in part (b). (f) Now calculate the ordinary least squares bo and b, for the five points. Label (g) Find the new residuals. Specifically, find e (h) Calculate the new sum of squared residuals (2 e?). them bo and bi. Y (box H, J, K, L, and M (i) Which line leads to the smaller sum of squared residuals? Does any line exist that leads to a smaller sum of squared residuals?
Answer #1

Anstwev Civen thal Name x-value y- Valve 5 rt (a) plot the Five Poin t and the centvoi d Desure yis the vertical axis and Vis the hori2onta ais ל The centvold of the dota base setisf0undgos Point are plotted as be0- Centoi J 2 3 니Draw a line that appearg f the Five point tue need bvau aline that passes through the centrial and is close to og many data point as posble Centra d 5 니 2 c) find the be ond bfor the line yo d vau in thep The bestit tine is astghith linçov eauatin find this equuation ue need tuwopoint ont are Ct ) and (315)(i) Find the residual isperifically calculate for J (2 is) ej For k (36) ek = 8-(6-13)=8-6 t325